OpenStudy (anonymous):

What is the solution set for |2x – 3| = 17?

OpenStudy (anonymous):

x = 10, -7

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

Yea he is

OpenStudy (valpey):

This is easy to confirm: 2*10-3 = 17 and 2*(-7) -3 = -17 absolute value of -17 = |-17| = 17